### v2

@@ 3.2 Model Training and Evaluation Metrics
The proposed cross-modal alignment and fusion network was implemented using the PyTorch framework (version 2.3.1) . All experiments were conducted on a workstation equipped with an NVIDIA GeForce RTX 3090 GPU (24 GB memory) running Ubuntu 22.04 LTS with CUDA 12.1 and Python 3.10.14.
@@ 3.4 Comparisons With Previous Methods
We benchmark CMAFNet against state-of-the-art object detection methods spanning three architectural paradigms: CNN-based single-stage detectors from the YOLO family (v5, v8, v10, v11), transformer-based end-to-end detectors built on the DETR framework (RT-DETR, Conditional-DETR, DETR, DINO), and a specialized small-object detector (FFCA-YOLO). Additionally, we compare against TinyDef-DETR, a domain-specific transformer tailored for transmission-line defect detection. For RGB-only baselines, we adopt results from our prior work (Shen et al., 2025), ensuring identical dataset splits and training configurations for fair comparison.
Table 3 presents the quantitative results on the TL-RGBD dataset. CMAFNet-x achieves the highest mAP50 of 0.322, surpassing the best-performing baseline DINO (0.224) by 9.8 percentage points. This superiority persists across all model scales: CMAFNet-n attains an mAP50 of 0.248 with only 4.9M parameters and 12.4 GFLOPs, outperforming all YOLO variants and matching larger transformer models. Such scaling behavior indicates that the performance gains stem primarily from cross-modal fusion rather than increased model capacity. For small-object detection—the core challenge in transmission-line imagery—CMAFNet-x achieves an APs of 0.125, improving upon DINO (0.085) by 4.0 percentage points and substantially exceeding YOLO variants (0.06–0.08). The mAP50:95 metric further confirms this advantage: CMAFNet-x reaches 0.147 versus 0.099 for DINO and 0.072–0.083 for leading YOLO configurations, demonstrating that cross-modal fusion enhances both detection recall and boundary localization. Although large objects constitute merely 0.02% of the TL-RGBD dataset (13 instances with area $\geq 96^{2}$ pixels), we report APl for completeness; notably, even the smallest variant CMAFNet-n achieves 0.369 on this subset, establishing a lower bound that larger configurations are expected to exceed. Category-specific metrics for jyzwh and dgss—the two most prevalent defect types in the TL-RGBD dataset—reveal pronounced improvements: CMAFNet-x achieves a recall of 0.584 and mAP50 of 0.517 for jyzwh defects (DINO: 0.437 and 0.420), and 0.463 recall with 0.465 mAP50 for dgss defects (DINO: 0.402 and 0.426). These results validate that depth modality provides discriminative geometric cues complementing RGB appearance features, particularly for defects with subtle visual signatures but distinctive three-dimensional profiles.
@@
YOLO v11n 300 2.6 6.3 0.638 0.135 0.136 0.064 0.060 0.103 263.2 0.353 0.274 0.324 0.245
YOLO v11s 300 9.4 21.3 0.516 0.166 0.155 0.072 0.074 0.100 125.0 0.381 0.314 0.344 0.291
YOLO v11s-P2 300 9.6 28.7 0.608 0.183 0.181 0.083 0.087 0.100 99.0 0.384 0.326 0.366 0.323
YOLO v11m 300 20.0 67.7 0.399 0.181 0.173 0.078 0.082 0.094 65.8 0.382 0.358 0.373 0.307
@@ 4 Conclusion
This paper presented CMAFNet, a Cross-Modal Alignment and Fusion Network that addresses the persistent challenges of RGB-D transmission-line defect detection through a principled “purify-then-fuse” paradigm. The proposed architecture systematically tackles three interrelated problems: intra-modal noise contamination arising from illumination artifacts and depth sensor limitations, cross-modal distribution heterogeneity that impedes effective feature alignment, and the absence of global contextual reasoning essential for disambiguating small-scale defects from visually similar background elements.
From a methodological perspective, CMAFNet introduces two complementary mechanisms that collectively enable robust multimodal fusion. The Semantic Recomposition Module performs dictionary-based feature purification by projecting representations into a learned codebook with controllable capacity, creating an information bottleneck that filters modality-specific noise while preserving defect-relevant patterns. The position-wise normalization within SRM transforms cross-modal alignment from implicit weight-based learning into explicit reconstruction-driven enhancement, ensuring that heterogeneous RGB and depth features attain compatible statistical distributions before fusion. The Contextual Semantic Integration Framework subsequently establishes global spatial dependencies through partial-channel attention, enabling the network to leverage structural priors—such as the regular arrangement of insulator strings and spatial relationships among transmission components—for semantic disambiguation. The partial-channel design, combined with the Adaptive Semantic Regulation Mechanism, prevents the detail erosion and over-smoothing that typically accompany full-channel global attention, thereby preserving the fine-grained discrimination essential for small-object detection.
Extensive experiments on the TL-RGBD dataset, where 94.5% of annotated instances qualify as small objects, demonstrate that CMAFNet achieves substantial improvements over state-of-the-art methods across all evaluation metrics. The full-scale model attains 32.2% mAP50 and 12.5% small-object AP, consistently outperforming RGB-based methods across all evaluation metrics. The lightweight configuration achieves 24.8% mAP50 at 228 FPS with only 4.9M parameters and 12.4 GFLOPs, outperforming all YOLO variants and matching larger transformer-based detectors at a fraction of their computational cost. Ablation studies further reveal a synergistic relationship between the proposed modules: while SRM and CSIF contribute 4.8% and 2.0% individual gains, their combined integration yields 13.7% total improvement, indicating non-linear complementary benefits that arise from addressing cross-modal distribution instability and local fusion limitations simultaneously. Modality ablation confirms that depth information provides complementary geometric cues—particularly for low-contrast defects and boundary disambiguation—that RGB appearance alone cannot capture, validating the necessity of cross-modal fusion for transmission-line inspection scenarios.
Several limitations warrant acknowledgment and suggest directions for future investigation. The current architecture assumes synchronized and spatially aligned RGB-D inputs, which may not hold under all deployment conditions involving diverse sensor configurations or dynamic acquisition platforms. The dictionary-based purification mechanism, while effective for the defect categories present in TL-RGBD, requires empirical tuning of codebook capacity when generalizing to domains with substantially different semantic complexity. Additionally, the performance gap between challenging categories—such as insulator contamination exhibiting subtle textural variations—and categories with distinctive geometric signatures indicates that adaptive fusion strategies conditioned on defect characteristics may yield further improvements. Future work will explore self-supervised pretraining on unlabeled multimodal inspection data to enhance feature representations, investigate dynamic fusion mechanisms that modulate cross-modal interaction based on input quality, and extend the framework to incorporate temporal consistency constraints for video-based inspection scenarios.

### v3

@@ 3.1 Datasets
For experimental evaluation, the TL-RGBD dataset was randomly divided into training, validation, and test subsets following an 8:1:1 split ratio. Specifically, 80% of the image pairs were used for model training, 10% for validation and hyperparameter selection, and the remaining 10% were reserved for final performance evaluation. The split was conducted at the image level to ensure that paired RGB and depth images always remain within the same subset. This partitioning strategy ensures sufficient training data while enabling unbiased validation and testing under consistent data distributions.
@@ 3.2 Model Training and Evaluation Metrics
The proposed cross-modal alignment and fusion network was implemented using the PyTorch framework (version 2.3.1) . All experiments were conducted on a workstation equipped with an NVIDIA GeForce RTX 3090 GPU (24 GB memory) running Ubuntu 22.04 LTS with CUDA 12.1 and Python 3.10.14.During training, a batch size of 16 was adopted for all experiments, balancing GPU memory constraints and optimization stability on the RTX 3090 platform.
@@ 3.4 Comparisons With Previous Methods
We benchmark CMAFNet against state-of-the-art object detection methods spanning three architectural paradigms: CNN-based single-stage detectors from the family (v5, v8, v10, v11), transformer-based end-to-end detectors built on the DETR framework (RT-DETR, Conditional-DETR, DETR, DINO), and a specialized small-object detector (FFCA-YOLO). Additionally, we compare against TinyDef-DETR, a domain-specific transformer tailored for transmission-line defect detection. For RGB-only baselines, we adopt results from our prior work (Shen et al., 2025), ensuring identical dataset splits and training configurations for fair comparison.
For all compared detectors, we followed their official training recipes, i.e., the default hyperparameter settings recommended in the corresponding open-source implementations or described in the original papers. To avoid unintentionally biasing the comparison through dataset-specific hyperparameter tuning, we did not modify algorithm-specific optimization strategies, including the optimizer type, learning-rate schedule, or assignment/matching rules. Only dataset-dependent adaptations were applied when strictly necessary, such as adjusting the number of categories and data loading paths.
Unless otherwise specified, all methods were trained and evaluated using the same input resolution and the same TL-RGBD train/validation/test split, ensuring a fair and controlled comparison across different architectural paradigms.
Table 3 presents the quantitative results on the TL-RGBD dataset. CMAFNet-x achieves the highest mAP50 of 0.322, surpassing the best-performing baseline DINO (0.224) by 9.8 percentage points. This superiority persists across all model scales: CMAFNet-n attains an mAP50 of 0.248 with only 4.9M parameters and 12.4 GFLOPs, outperforming all YOLO variants and matching larger transformer models. Such scaling behavior indicates that the performance gains stem primarily from cross-modal fusion rather than increased model capacity. For small-object detection—the core challenge in transmission-line imagery—CMAFNet-x achieves an APs of 0.125, improving upon DINO (0.085) by 4.0 percentage points and substantially exceeding YOLO variants (0.06–0.08). The mAP50:95 metric further confirms this advantage: CMAFNet-x reaches 0.147 versus 0.099 for DINO and 0.072–0.083 for leading YOLO configurations, demonstrating that cross-modal fusion enhances both detection recall and boundary localization. Although large objects constitute merely 0.02% of the TL-RGBD dataset (13 instances with area $\geq 96^{2}$ pixels), we report APl for completeness; notably, even the smallest variant CMAFNet-n achieves 0.369 on this subset, providing a conservative reference point that larger configurations are expected to surpass. Category-specific metrics for jyzwh and dgss—the two most prevalent defect types in the TL-RGBD dataset—reveal pronounced improvements: CMAFNet-x achieves a recall of 0.584 and mAP50 of 0.517 for jyzwh defects (DINO: 0.437 and 0.420), and 0.463 recall with 0.465 mAP50 for dgss defects (DINO: 0.402 and 0.426). These results validate that depth modality provides discriminative geometric cues complementing RGB appearance features, particularly for defects with subtle visual signatures but distinctive three-dimensional profiles.
@@
YOLO 11n 300 2.6 6.3 0.638 0.135 0.136 0.064 0.060 0.103 263.2 0.353 0.274 0.324 0.245
YOLO 11s 300 9.4 21.3 0.516 0.166 0.155 0.072 0.074 0.100 125.0 0.381 0.314 0.344 0.291
YOLO 11s-P2 300 9.6 28.7 0.608 0.183 0.181 0.083 0.087 0.100 99.0 0.384 0.326 0.366 0.323
YOLO 11m 300 20.0 67.7 0.399 0.181 0.173 0.078 0.082 0.094 65.8 0.382 0.358 0.373 0.307
@@ 4 Conclusion
This paper has presented CMAFNet, a cross-modal alignment and fusion network that follows a principled purify-then-fuse paradigm for RGB-D transmission-line defect detection. The central insight motivating this design is that multimodal fusion operating on noisy, statistically misaligned representations inherits and amplifies their deficiencies rather than compensating for them—a problem that is especially acute when the vast majority of targets are small objects occupying only a few dozen pixels. CMAFNet addresses this challenge through two complementary modules. The Semantic Recomposition Module projects features through a learned bottleneck with position-wise normalization, suppressing modality-specific artifacts—illumination variation and specular noise in RGB, quantization errors and edge bleeding in depth—while narrowing the inter-modal distribution gap before integration. The Contextual Semantic Integration Framework then establishes global spatial dependencies through partial-channel attention at the deepest fusion level, enabling the network to exploit structural priors such as the regular arrangement of insulator strings and the spatial relationships between fittings and conductors for disambiguating small defects from visually similar background elements, without the detail erosion that full-channel attention would impose on fine-grained localization.
Extensive experiments on the TL-RGBD benchmark, where 94.5% of annotated instances qualify as small objects, validate the effectiveness of the proposed approach. The full-scale variant CMAFNet-x achieves 32.2% mAP50 and 12.5% APs, surpassing the strongest baseline by 9.8 and 4.0 percentage points respectively, while the lightweight configuration CMAFNet-n attains 24.8% mAP50 at 228 FPS with only 4.9M parameters—meeting the throughput requirements of real-time UAV-based inspection. Ablation studies reveal a synergistic interaction between SRM and CSIF: their combined integration yields a 13.7% relative improvement in mAP50, substantially exceeding the sum of their individual contributions, confirming that semantic purification and global context modeling address distinct yet mutually reinforcing bottlenecks. Modality ablation further demonstrates that depth provides complementary geometric cues—particularly for boundary disambiguation and low-contrast defect localization—that RGB appearance alone cannot capture, with the fused configuration achieving a 26.4% relative gain in mAP50 over the RGB-only baseline. Feature heatmap visualizations corroborate these quantitative findings, showing that the proposed modules transform diffuse, background-spread activations into spatially concentrated patterns tightly aligned with defect locations.
Despite these encouraging results, several limitations warrant discussion. The current architecture restricts cross-modal fusion to $P_{4}$ and $P_{5}$, deliberately excluding depth at $P_{3}$ to avoid injecting high-noise sensor signals into the feature map most critical for small-object localization. While this design choice proves effective under the noise characteristics of the Intel RealSense D435i sensor used in TL-RGBD, it forecloses the possibility of leveraging fine-grained geometric detail at higher resolutions—a limitation that may become significant as depth sensing hardware improves in precision and spatial fidelity. A noise-adaptive gating mechanism that dynamically modulates depth contribution at each pyramid level according to estimated signal quality would offer a more flexible alternative to the current fixed exclusion. Furthermore, the absolute detection performance, though substantially improved over prior methods, remains modest in terms of mAP50:95, reflecting the intrinsic difficulty of achieving precise localization for targets that span fewer than $32\times 32$ pixels. Advancing small-object AP beyond the current level will likely require joint innovation in both feature representation and label assignment strategies that account for the spatial uncertainty inherent in bounding-box supervision at this scale.
From a broader perspective, the purify-then-fuse paradigm validated here is not restricted to the RGB-depth pairing investigated in this work. Transmission-line inspection increasingly involves heterogeneous sensor suites—including thermal infrared cameras for hotspot detection and LiDAR for three-dimensional structural modeling—whose integration faces analogous challenges of distributional mismatch and modality-specific noise. Extending the proposed framework to accommodate additional modalities, potentially through modality-agnostic purification modules that share a common bottleneck architecture while learning modality-specific projection parameters, constitutes a natural direction for future research. Equally important is the evaluation of CMAFNet on diverse infrastructure inspection benchmarks beyond transmission lines, where the combination of small-object prevalence and complex backgrounds similarly constrains single-modal approaches. Finally, while the lightweight CMAFNet-n variant already demonstrates real-time capability on GPU hardware, deploying the model on resource-constrained UAV platforms with limited power and memory budgets will require further exploration of model compression techniques—including knowledge distillation from the full-scale variant and structured pruning guided by modality-specific sensitivity analysis—to bridge the gap between laboratory throughput and field-deployable efficiency.